OpenStudy (anonymous):

Among all pairs of numbers whose difference is 10, find a pair whose product is as small as possible. What is the minimum product? a. the pairs of numbers whose difference is 10 and whose product is as small as possible is___ b. The minimum product is___

OpenStudy (anonymous):

\((10,0)\) Answer b yourself.

OpenStudy (anonymous):

hmm what about -5 and 5?

OpenStudy (anonymous):

Hm. That may very well be the best pair. Do you have a proof?

OpenStudy (anonymous):

ok that was wrong but idea was right

OpenStudy (anonymous):

try with \(x\) and \(x-10\)

OpenStudy (anonymous):

Hm. So \(x-y=10\), \(y=x-10\), and \(xy=x(x-10)=x^2-10x\). The minima of \(x^2-10x\) is at \(x=5\), so \(x=5\), \(y=5-10=-5\). Nice job.